OpenStudy (anonymous):

Anyone willing to help me find this one out? A gym offers regular memberships for $80 per month and off-peak memberships for $60 per month. Last month, the gym sold a total of 420 memberships for a total of $31,100. How many memberships sold were regular memberships? 125 140 235 295

OpenStudy (anonymous):

x + y = 420 80x + 60y = 31,100 Solve the system of equations

OpenStudy (anonymous):

81x+61y=3,520. do i go on from there?

OpenStudy (anonymous):

if so, do i solve for x or for y?

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (anonymous):

@Hero @satellite73

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

@Hero